# Multitask-based Evaluation of Open-Source LLM on Software Vulnerability

### v2

@@ 4.3 RQ-3: Evaluating Vulnerability Location of LLMs
As for a specific vulnerable function, it may contain one or several vulnerable lines of code ($Lines_{ground}$), and LLM may also predict one or several potential ones ($Lines_{predict}$). We compare $Lines_{predict}$ to $Lines_{ground}$ to check whether the line index $L_{i}\in Lines_{predict}$ belongs to $Lines_{ground}$ and we treat it predict correctly for the line $L_{i}$ if it belongs to, otherwise it predicts incorrectly. We also use the $Lines_{both}$ to represent the intersection of $Lines_{predict}$ to $Lines_{ground}$.
To better evaluate the vulnerability location performance of LLM on a specific vulnerable function, we give the following definitions:
  • Hit@Acc means the effectiveness of LLM and equals 1 if LLM correctly predicts at least one line of vulnerable line, other it equals 0.
[…]
For example, for a given vulnerable function, it totally has six vulnerable lines “[2, 3, 5, 9, 14,23]”, and LLM gives out its prediction with 10 potential lines “[1, 3, 5, 11, 15, 16, 17, 21, 22, 23]”. Then, we know that $Lines_{ground}$ equals “[2, 3, 5, 9, 14, 23]”, $Lines_{predict}$ equals “[1, 3, 5, 11, 15, 16, 17, 21, 22, 23]” and $Lines_{both}$ equals “[3, 5, 23]”. According to these values, we obtain that $Precision=\frac{3}{10}$, $Recall=\frac{3}{6}$, and $Hit@Acc=1$.
TABLE IX: The comparison of LLMs on software vulnerability location
Methods Precision Recall Hit@Acc
@@
DeepSeek-Coder 6.7B 0.121 0.186 0.307
CodeLlama 7B 0.097 0.137 0.304
StarCoder 7B 0.113 0.175 0.284
WizardCoder 7B 0.111 0.122 0.255
Mistral 0.133 0.163 0.320
Phi-2 0.081 0.177 0.276
@@
DeepSeek-Coder 6.7B 0.085 0.314 0.472
DeepSeek-Coder 33B 0.094 0.393 0.508
CodeLlama 7B 0.066 0.228 0.335
CodeLlama 34B 0.082 0.231 0.358
StarCoder 7B 0.093 0.341 0.469
StarCoder 15.5B 0.085 0.499 0.662
WizardCoder 7B 0.078 0.188 0.338
WizardCoder 34B 0.086 0.204 0.327
Mistral 0.068 0.060 0.155
Phi-2 0.102 0.063 0.211
TABLE X: The comparison of software vulnerability location performance for Top-10 CWE Types between LLMs
CWE Type # Vul. DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2 DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2
Precision of Fine-Tuning Setting Precision of Few-Shot Setting
CWE-119 128 0.121 0.065 0.125 0.103 0.100 0.056 0.084 0.043 0.099 0.052 0.043 0.105
CWE-20 80 0.154 0.141 0.198 0.190 0.178 0.088 0.075 0.093 0.097 0.121 0.073 0.054
CWE-264 64 0.083 0.065 0.112 0.059 0.090 0.081 0.064 0.062 0.072 0.058 0.081 0.091
CWE-399 35 0.108 0.194 0.128 0.116 0.173 0.055 0.066 0.050 0.131 0.058 0.036 0.152
CWE-125 29 0.094 0.127 0.048 0.055 0.090 0.112 0.087 0.077 0.090 0.066 0.041 0.065
CWE-200 27 0.088 0.082 0.041 0.077 0.120 0.062 0.068 0.072 0.078 0.066 0.092 0.127
CWE-189 21 0.048 0.170 0.035 0.024 0.231 0.294 0.125 0.083 0.069 0.043 0.041 0.185
CWE-362 16 0.037 0.069 0.027 0.088 0.138 0.077 0.032 0.034 0.089 0.030 0.033 0.167
CWE-416 12 0.200 0.333 0.031 0.286 0.080 0.034 0.203 0.116 0.134 0.185 0.000 0.190
CWE-476 11 0.020 0.053 0.029 0.111 0.071 0.000 0.120 0.099 0.129 0.256 0.200 0.071
CWE Type # Vul. DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2 DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2
Recall of Fine-Tuning Setting Recall of Few-Shot Setting
CWE-119 128 0.184 0.128 0.234 0.128 0.163 0.153 0.356 0.163 0.366 0.147 0.038 0.078
CWE-20 80 0.144 0.106 0.116 0.106 0.106 0.069 0.156 0.153 0.275 0.169 0.041 0.025
CWE-264 64 0.138 0.114 0.263 0.114 0.120 0.293 0.341 0.305 0.341 0.198 0.096 0.096
CWE-399 35 0.317 0.256 0.256 0.195 0.366 0.244 0.293 0.220 0.463 0.146 0.024 0.085
CWE-125 29 0.208 0.178 0.089 0.059 0.099 0.297 0.356 0.267 0.327 0.119 0.030 0.030
CWE-200 27 0.262 0.123 0.062 0.092 0.185 0.215 0.400 0.354 0.415 0.323 0.169 0.108
CWE-189 21 0.077 0.205 0.103 0.026 0.615 0.128 0.615 0.359 0.436 0.154 0.103 0.128
CWE-362 16 0.125 0.250 0.125 0.375 1.000 0.500 0.125 0.125 0.625 0.125 0.125 0.125
CWE-416 12 0.356 0.289 0.044 0.311 0.044 0.022 0.622 0.378 0.422 0.378 0.000 0.089
CWE-476 11 0.071 0.286 0.071 0.286 0.143 0.000 0.786 0.786 0.786 0.714 0.286 0.071
CWE Type # Vul. DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2 DeepSeek-Coder CodeLlama StarCoder WizardCoder Mistral Phi-2
Hit@Acc of Fine-Tuning Setting Hit@Acc of Few-Shot Setting
CWE-119 128 0.326 0.292 0.270 0.191 0.270 0.292 0.416 0.270 0.438 0.303 0.101 0.247
CWE-20 80 0.404 0.426 0.426 0.404 0.532 0.277 0.489 0.362 0.596 0.383 0.149 0.149
CWE-264 64 0.172 0.241 0.414 0.276 0.276 0.259 0.397 0.276 0.379 0.328 0.121 0.224
CWE-399 35 0.476 0.286 0.190 0.238 0.524 0.333 0.571 0.286 0.476 0.238 0.095 0.238
CWE-125 29 0.500 0.364 0.318 0.227 0.273 0.318 0.545 0.545 0.591 0.273 0.136 0.136
CWE-200 27 0.385 0.231 0.154 0.192 0.385 0.269 0.538 0.500 0.577 0.577 0.346 0.269
CWE-189 21 0.250 0.333 0.250 0.083 0.417 0.333 0.667 0.333 0.667 0.500 0.250 0.333
CWE-362 16 0.333 0.667 0.333 0.333 1.000 0.333 0.333 0.333 0.667 0.333 0.333 0.333
CWE-416 12 0.200 0.400 0.200 0.400 0.200 0.100 0.600 0.300 0.500 0.300 0.000 0.400
CWE-476 11 0.143 0.429 0.143 0.429 0.286 0.000 0.571 0.571 0.571 0.571 0.571 0.143
Results. Table IX presents the average performance of vulnerability location for LLMs, while Table X provides a comparison of software vulnerability location performance among the Top-10 CWE Types across different LLMs. Both Table IX and Table X provide a comparative analysis of the impact of fine-tuning and few-shot settings on LLMs. Based on these tables, we can achieve the following observations:
(1) The few-shot setting reveals limitations in LLMs, and after fine-tuning, LLMs become more cautious. From the perspective of the Recall and Hit@Acc metrics, it appears that the vulnerability location capability of the LLMs after fine-tuning has actually weakened. For example, as shown in Table IX, the Recall range has changed from 0.063-0.341 to 0.122-0.186, and the Hit@Acc range has changed from 0.155-0.472 to 0.255-0.320. From Table X, it can be observed that across different CWE types, the best performance in terms of Recall and Hit@Acc metrics often occurs in the few-shot setting. This contradicts our previous experience, as it seems that LLMs perform worse after learning. To address this issue, we conducted a detailed analysis of the output results of the LLMs under fine-tuning and few-shot settings. We found that LLMs have limitations, and in the few-shot setting, they tend to output more vulnerable lines, even if these lines do not contain vulnerabilities. Let’s take StarCoder as an example. Fig. 5 depicts a vulnerability code snippet from the Big-Vul dataset, with the vulnerability behavior occurring in lines 3 and 4. However, in the few-shot setting, StarCoder tends to output more vulnerability lines, such as “[1, 2, 3, 4, 5, 6, 7, 8, 9]”, whereas after fine-tuning, StarCoder becomes more cautious and only outputs “[4]”. We observed a similar pattern in other LLMs, and we believe this is the reason for the higher Recall and Hit@Acc values in the few-shot setting. Therefore, it appears that the excellent performance of LLMs in few-shot setting is attributed to their inclination to output more potential vulnerability lines in those scenarios, even when such lines do not contain actual vulnerabilities.
Figure 5: An example to demonstrate the limitations of StarCoder in vulnerability location
(2) Mistral is a promising LLM for vulnerability location task. Overall, after fine-tuning, LLMs show improved Precision, indicating their ability to learn and perform better on vulnerability location tasks. Surprisingly, Mistral, being a text LLM, exhibits improvements across all metrics after fine-tuning, especially with Precision reaching the best (i.e., 0.133), surpassing all other LLMs. This suggests that Mistral possesses significant potential to learn how to effectively locate vulnerabilities. If Mistral could be trained using pre-training tasks similar to those used for code LLMs, it might become an exceptionally powerful model for software vulnerability location.
Finding-6. Few-shot settings expose LLM limitations, but fine-tuning enhances caution. Finding-7. Mistral’s significant improvement after fine-tuning showcases its potential.
@@ 4.4 RQ-4: Evaluating Vulnerability Description of LLMs
Experimental Setting. We instruct LLMs with a designated task description, guiding it to perform the role of a vulnerability descriptor. Table XII illustrates an example of our approach to evaluating LLMs’ proficiency in conducting vulnerability descriptions.
@@
TABLE XI: The comparison of LLMs on software vulnerability description
@@
DeepSeek-Coder 6.7B 0.296 0.125 0.281
CodeLlama 7B 0.505 0.360 0.493
StarCoder 7B 0.532 0.397 0.522
WizardCoder 7B 0.514 0.373 0.502
Mistral 0.536 0.409 0.528
Phi-2 0.308 0.110 0.288
@@
DeepSeek-Coder 6.7B 0.237 0.078 0.218
DeepSeek-Coder 33B 0.232 0.066 0.216
CodeLlama 7B 0.232 0.088 0.219
CodeLlama 34B 0.247 0.072 0.227
StarCoder 7B 0.229 0.081 0.215
StarCoder 15.5B 0.238 0.080 0.220
WizardCoder 7B 0.218 0.061 0.202
WizardCoder 34B 0.254 0.098 0.240
Mistral 0.264 0.100 0.249
Phi-2 0.253 0.104 0.241
TABLE XII: An example of CodeLlama generating vulnerability description for vulnerable code
Source Code An example of a C code snippet with vulnerabilities. CVE ID: CVE-2017-0637.
Response From CodeLlama A remote denial of service vulnerability in libhevc in Mediaserver could enable an attacker to use a specially crafted file to cause a device hang or reboot. This issue is rated as High severity due to the possibility of remote denial of service. Product: Android. Versions: 5.0.2, 5.1.1, 6.0, 6.0.1, 7.0, 7.1.1, 7.1.2. Android ID: A-34819017. Android ID: A-34819017.
CVE Description A remote code execution vulnerability in libhevc in Mediaserver could enable an attacker using a specially crafted file to cause memory corruption during media file and data processing. This issue is rated as Critical due to the possibility of remote code execution within the context of the Mediaserver process.Product: Android. Versions: 5.0.2, 5.1.1, 6.0, 6.0.1, 7.0, 7.1.1, 7.1.2. Android ID: A-34064500.
Results. Table XI represents the vulnerability description capabilities of LLMs on difference settings. According to the results, we can obtain the following observations: (1) Fine-tuning can significantly enhance the performance of LLMs in vulnerability descriptions, especially in the case of CodeLlama, StarCoder, WizardCoder, and Mistral. After fine-tuning, there is a several-fold improvement in ROUGE-1, ROUGE-2, and ROUGE-L metrics. This suggests that these LLMs possess strong learning capabilities and can extract more gains from historical data. (2) The low ROUGE-2 scores indicate that DeepSeek-Coder and Phi-2 have limited ability to generate accurate and relevant high-order n-grams (pairs of consecutive words) in vulnerability descriptions, indicating potential issues in capturing specific and detailed information.
@@
Case Study. To demonstrate the effectiveness of LLMs in generating vulnerability descriptions, we present an example of a vulnerability (CWE-119) described by CodeLlama, as shown in Table XII. This example represents a vulnerability within the Linux project, categorized as CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer Vulnerability). It is noteworthy that even when provided with only the code of the vulnerability, CodeLlama produces text highly similar to the CVE description (highlighted in orange), indicating that pre-trained LLMs on extensive text and code are capable of comprehending the essence and crucial features of vulnerabilities and expressing this information in natural language.
Finding-8: CodeLlama, StarCoder, WizardCoder, and Mistral possess strong learning capabilities and can extract more gains from historical data.
5 Threats to Validate

### v3

@@ 4.3 RQ-3: Evaluating Vulnerability Location of LLMs
For the fine-tuning setting of LLMs and pre-trained LMs, we treat the vulnerability location task as a binary classification problem, determining whether each line of code is vulnerable or not. For the few-shot setting, a specific vulnerable function may contain one or several vulnerable lines, and LLM may also predict one or several potential vulnerable lines ($Lines_{predict}$). We convert $Lines_{predict}$ into a binary classification format. For example, if a given vulnerable function consists of five lines and contains two vulnerable lines [2, 3], and the LLM predicts one potential vulnerable line [2], we convert this to a binary classification format as [0, 0, 1, 0, 0] for ease of computation. To better evaluate the vulnerability location performance of LLM on a specific vulnerable function, we consider five widely used performance measures (i.e., Precision, Recall, F1-score, Accuracy, and FPR).
In addition to pre-trained LMs, we also consider the following four SOTA baselines: Devign [38], Reveal [47], IVDetect [56], and LineVul [39]. For the graph-based approaches (i.e., Devign, Reveal, and IVDetect), we use GNNExplainer [78, 79] for vulnerability location. We compare the performance of LLMs and these baselines using Top-k Accuracy, as employed in previous works [39, 79].
TABLE X: The comparison between LLMs and six baselines on software vulnerability location (RQ3)
Methods F1-score Recall Precision Accuracy FPR
CodeBERT 0.470 0.514 0.433 0.879 0.078
GraphCodeBERT 0.483 0.477 0.489 0.893 0.058
UniXcoder 0.460 0.384 0.575 0.908 0.032
PLBART 0.436 0.416 0.458 0.886 0.058
CodeT5 0.493 0.408 0.623 0.914 0.028
CodeT5+ 0.303 0.207 0.565 0.902 0.018
@@
DeepSeek-Coder 6.7B 0.437 0.332 0.640 0.912 0.021
CodeLlama 7B 0.504 0.396 0.691 0.919 0.021
StarCoder 7B 0.245 0.169 0.443 0.893 0.024
WizardCoder 7B 0.520 0.427 0.664 0.918 0.025
Mistral 0.314 0.384 0.266 0.827 0.122
Phi-2 0.458 0.361 0.629 0.912 0.025
@@
DeepSeek-Coder 6.7B 0.111 0.111 0.112 0.852 0.081
DeepSeek-Coder 33B 0.110 0.112 0.108 0.849 0.084
CodeLlama 7B 0.082 0.063 0.116 0.882 0.043
CodeLlama 34B 0.115 0.090 0.158 0.884 0.044
StarCoder 7B 0.088 0.066 0.134 0.887 0.039
StarCoder 15.5B 0.095 0.078 0.120 0.876 0.052
WizardCoder 7B 0.082 0.063 0.120 0.884 0.042
WizardCoder 34B 0.096 0.072 0.145 0.887 0.039
Mistral 0.086 0.065 0.127 0.885 0.040
Phi-2 0.073 0.053 0.116 0.885 0.037
Results. Table X presents the overall performance of vulnerability location between LLMs and seven baselines. Based on this table, we can achieve the following observations: (1) Fine-tuning can greatly enhance the vulnerability location capabilities of LLMs. For example, after fine-tuning, CodeLlama 7B’s F1-score increases from 0.082 to 0.504, recall increases from 0.063 to 0.396, precision increases from 0.116 to 0.691, accuracy increases from 0.882 to 0.919, and FPR decreases from 0.043 to 0.021. (2) Code-related LLMs often outperform pre-trained LMs in terms of F1-score, precision, and FPR. For example, CodeLlama 7B outperforms the pre-trained LMs, which are averaged over six different models, by 14.3%, 31.9%, and 54.3% in terms of F1-score, precision, and FPR, respectively.
The Top-k Accuracy results of the interpreters are shown in Fig. 7, where the x-axis represents k and the y-axis represents Top-k Accuracy (%). For comparison, we average the results of all LLMs and pre-trained LMs. We can observe that the performance of these graph-based methods does not show significant differences but is considerably weaker than that of pre-trained LMs, LLMs, and LineVul. Although pre-trained LMs achieve the highest accuracy at k=20, the difference between them and LLMs is not substantial.
TABLE XI: An example of generating vulnerability description for vulnerable code (RQ4)
Source Code An example of a C code snippet with vulnerabilities. CVE ID: CVE-2016-0841.
Response From CodeT5+ libmedia in mediaserver in Android 4.x before 4.4.4, 5.0.x before 5.0.2, 5.1.x before 5.1.1, and 6.x before 2016-08-01 has certain incorrect declarations, which allows remote attackers to execute arbitrary code or cause a denial of service (NULL pointer dereference or memory corruption) via a crafted media file, aka internal bug 28166152.
Response From CodeLlama media/libmedia/MediaMetadataRetriever.cpp in mediaserver in Android 4.x before 4.4.4, 5.0.x before 5.0.2, 5.1.x before 5.1.1, and 6.x before 2016-04-01 does not validate the data size, which allows attackers to execute arbitrary code or cause a denial of service (memory corruption) via a crafted media file, aka internal bug 28799341. Specifically, the constructor calls the native function createMetadataRetriever in libstagefright, and later calls the getFrameAtTime function, aka internal bug 29244311. The getFrameAtTime call triggers a memory corruption warning, which allows an attacker to execute arbitrary code via a crafted video file, aka internal …
CVE Description media/libmedia/mediametadataretriever.cpp in mediaserver in Android 4.x before 4.4.4, 5.0.x before 5.0.2, 5.1.x before 5.1.1, and 6.x before 2016-04-01 mishandles cleared service binders, which allows remote attackers to execute arbitrary code or cause a denial of service (memory corruption) via a crafted media file, aka internal bug 26040840.
Figure 7: The interpretation results for different vulnerability locators with k from 2 to 20 (RQ3)
Case Study. We find that LLMs under few-shot setting have limitations, and in some exceptional cases, they tend to output more vulnerable lines, even if these lines do not contain vulnerabilities. We take StarCoder as an example, Fig. 8 depicts a vulnerability code snippet from the Big-Vul dataset, with the vulnerability behavior occurring in lines 3 and 4. However, StarCoder tends to output more vulnerability lines, such as “[1, 2, 3, 4, 5, 6, 7, 8, 9]”, whereas after fine-tuning, StarCoder becomes more cautious and only predicts “[4]”. Note that we convert the model’s predictions into a specific format, i.e., transform “[0, 0, 0, 1, 0, 0, 0, 0, 0]” to “[4]”.
Figure 8: An example to demonstrate the limitations of StarCoder in vulnerability location (RQ3)
Finding-6. Few-shot setting exposes LLM’s limitations, and fine-tuning can greatly enhance the vulnerability location capabilities of LLMs. Finding-7. Fine-tuning code-related LLMs as vulnerability locators is beneficial, as they can outperform pre-trained LMs in terms of F1-score, precision, and FPR.
@@ 4.4 RQ-4: Evaluating Vulnerability Description of LLMs
Experimental Setting. We instruct LLMs with a designated task description, guiding them to perform the role of a vulnerability descriptor. Table XI illustrates an example of our approach to evaluating LLMs’ proficiency in conducting vulnerability descriptions.
@@
Results. Table XII represents the vulnerability description capabilities of LLMs and six baselines. According to the results, we can obtain the following observations: (1) LLMs exhibit significantly weaker performance in generating vulnerability descriptions compared to pre-trained LMs. For instance, pre-trained LMs achieve an average performance of 0.600, 0.487, and 0.591 on ROUGE-1, ROUGE-2, and ROUGE-L, respectively, whereas fine-tuned LLMs only achieve an average of 0.406, 0.301, and 0.400 on the same metrics. (2) Fine-tuning can significantly enhance the performance of LLMs in vulnerability descriptions. After fine-tuning, there is a several-fold improvement in ROUGE-1, ROUGE-2, and ROUGE-L. This suggests that these LLMs possess strong learning capabilities and can extract more gains from historical data. (3) The low ROUGE-2 scores indicate that Phi-2 has a limited ability to generate accurate and relevant high-order n-grams (pairs of consecutive words) in vulnerability descriptions, indicating potential issues in capturing specific and detailed information.
TABLE XII: The comparison of LLMs on software vulnerability description (RQ4)
@@
CodeBERT 0.511 0.376 0.501
GraphCodeBERT 0.538 0.406 0.528
UniXcoder 0.658 0.558 0.650
PLBART 0.447 0.313 0.437
CodeT5 0.700 0.604 0.693
CodeT5+ 0.747 0.668 0.740
@@
DeepSeek-Coder 6.7B 0.434 0.325 0.425
CodeLlama 7B 0.392 0.292 0.387
StarCoder 7B 0.420 0.321 0.416
WizardCoder 7B 0.425 0.327 0.419
Mistral 0.453 0.347 0.448
Phi-2 0.313 0.196 0.305
@@
DeepSeek-Coder 6.7B 0.230 0.073 0.215
DeepSeek-Coder 33B 0.219 0.066 0.203
CodeLlama 7B 0.221 0.070 0.205
CodeLlama 34B 0.258 0.094 0.242
StarCoder 7B 0.243 0.084 0.229
StarCoder 15.5B 0.255 0.089 0.241
WizardCoder 7B 0.230 0.066 0.211
WizardCoder 34B 0.276 0.111 0.261
Mistral 0.290 0.095 0.267
Phi-2 0.210 0.056 0.194
@@
Case Study. To demonstrate the capability of pre-trained LMs and LLMs in generating vulnerability descriptions, we present an example of a vulnerability (CWE-119) described by CodeT5+ and CodeLlama, as shown in Table XI. This example represents a vulnerability within the Linux project, categorized as CWE-119 (Improper Restriction of Operations within the Bounds of a Memory Buffer Vulnerability). It is noteworthy that even when provided with only the code of the vulnerability, CodeT5+ produces text highly similar to the CVE description (highlighted in orange), indicating that pre-trained LMs are capable of comprehending the essence and crucial features of vulnerabilities and expressing this information in natural language.
Additionally, we find that CodeLlama’s response is very similar to the CVE description, but with many additional details. We hypothesize that the poor performance of LLMs is not due to their inability to generate appropriate vulnerability descriptions, but rather because they tend to output tokens endlessly, even when they should stop. In contrast, pre-trained LMs typically stop at the appropriate points.
To further analyze this, we investigate the vulnerability description capabilities of LLMs after mitigating this issue. Using CodeLlama as an example, we randomly select 100 examples from the testing set and manually determine where the descriptions should terminate, trimming CodeLlama’s output accordingly. We then calculate the ROUGE metrics for the trimmed outputs and compare them with the original results and those of CodeT5+. The final results are presented in Table XIII, we find that after trimming, the ROUGE-1, ROUGE-2, and ROUGE-L scores for CodeLlama significantly improved, even nearing those of CodeT5+. This confirms our hypothesis that LLMs actually possess strong vulnerability description capabilities, but their performance is hindered by the tendency to output excessively.
TABLE XIII: The comparison of CodeT5+, CodeLlama, and CodeLlama-Trim on selected examples (RQ4)
Methods ROUGE-1 ROUGE-2 ROUGE-L
CodeT5+ 0.730 0.644 0.722
CodeLlama 0.366 0.266 0.360
CodeLlama-Trim 0.625 0.523 0.616
Finding-8. LLMs exhibit significantly weaker performance in generating vulnerability descriptions compared to pre-trained LMs. Therefore, fine-tuning pre-trained LMs for vulnerability detection is recommended.
5 Discussion
This section discusses open questions regarding the performance differences observed, the impact of input sequence length, and potential threats to the validity of our results.
5.1 Analysis of Performance Difference
In RQ1, for LineVul, there is a huge difference between the results obtained in this paper (i.e., 0.272 of F1-score) and the ones reported in original work (i.e., 0.910 of F1-score). To ensure a fair comparison, we first check the correctness of our LineVul reproduction by re-conducting the corresponding experiments using the original dataset provided by LineVul’s official source and we obtain similar results. Then, we inspect each step of the data preprocessing process, as outlined in Section 3.1. In particular, this process involves three pre-processing in total: removing blank lines, removing comments, and trimming leading and trailing spaces from lines. We pre-process the original dataset of LineVul, re-train, and test the model under the same parameter settings. The results are shown in Table XIV and we obtain the following conclusions:
  • Our reproduced LineVul performs closely to the original one.
  • Removing blank lines and comments does not significantly affect LineVul’s results.
  • Trimming leading and trailing spaces from lines causes a drastic decrease in LineVul’s performance.
Generally, for C/C++ source code, we know that removing leading and trailing spaces does not affect the code’s semantics. Thus, to verify whether it is general to other transformer-based models, we conduct another experiment on UniXcoder (another famous and widely used transformer-based pre-trained model) by adopting the same filtering operations. The results are presented in the right part of Table XIV. Table XIV shows that the UniXcoder’s performance closely resembled LineVul’s before the third step of processing. However, after pre-processing, UniXcoder’s performance similarly plummeted. Thus, we believe that such types of operation will have side impacts on transformer-based models since these methods pay attention to each token, though these tokens have no semantic meaning in the context of source code. Based on this observation, we believe that the vulnerability detection effectiveness of LineVul after space removal is correct and the performance results are reasonable.
TABLE XIV: The reproduced results for LineVul and UniXcoder
Datasets LineVul UniXcoder
F1-score Accuracy Recall Precision F1-score Accuracy Recall Precision
Original dataset 0.90 0.95 0.86 0.95 0.86 0.98 0.82 0.90
Remove empty lines 0.85 0.98 0.79 0.93 0.85 0.98 0.82 0.88
Remove comments 0.86 0.99 0.81 0.93 0.85 0.98 0.81 0.90
Remove spaces 0.40 0.94 0.37 0.45 0.26 0.95 0.16 0.68
5.2 Analysis of Input Sequence Length
In RQ2, we find that after adding key information, the performance of StarCoder and Phi-2 in vulnerability assessment actually weakened. We hypothesize that these LLMs may not excel in handling excessively long text sequences. Therefore, adding key information, which results in an increase in the number of input tokens, leads to a decline in performance. In this section, we aim to analyze the performance of StarCoder and Phi-2 with respect to input sequence length to determine whether there is a performance decline as the input length increases. As shown in Fig. 9, the horizontal axis represents the token length of the input sequence, and the vertical axis represents the F1-score of vulnerability assessment. We categorize the input token lengths into 0-128, 128-256, 256-512, 512-1024, and 1024+ (e.g., an input token length of 64 falls into the 0-128 range), and evaluate the vulnerability assessment performance of StarCoder and Phi-2 for each category. According to Fig. 9, we observe that as the input length increases, the F1-scores of both LLMs gradually decrease, revealing their significant limitations in assessing long sequences of vulnerable code. Therefore, in practical applications requiring the assessment of long sequences of vulnerable code, we may need to consider alternative optimization strategies or model choices to ensure accuracy and reliability.
Figure 9: The variation of F1-scores for vulnerability assessment with respect to input sequence length
5.3 Threats to Validity